# Optimizing Optimizers for Fast Gradient-Based Learning

### v1

@@ Appendix D Automating validation-aware optimizer tuning
Proposition 10 (Validation optimality in power).
@@
Figure 5: Demonstration of effectiveness of validation-aware design of gradient-based optimizers. The validation-aware optimizers achieve the highest test accuracy among all optimizers. The SGD+M optimizer is trained on the CIFAR-100 dataset (krizhevsky2009learning) with ResNet-18 (he2016deep).
@@
Although, from Proposition 10 and from our experiment, we know that validation-aware optimal optimizer gives instantaneously favorable choice over the training-only optimizer, a deeper analysis is required to understand and measure the exact effect of validation-awareness. We leave this as future work and resort to empirical evaluation for now. This concludes our theoretical and empirical demonstration of automatic, validation-aware design of gradient-based optimizers.
@@ Appendix E Reverse engineering common optimizers
In the main manuscript, we have provided the optimal forms of two of the most widely used optimizers: SGD with momentum and Adam (kingma2015adam), and postponed the extension to other optimizers to this appendix, leaving only Table 7 with the full list of optimizers in the main text. This section completes the reverse engineering of various optimizers under our greedy paradigm.
Our results include, but are not limited to, the following optimizers: SGD, Adam/AdamW (kingma2015adam; loshchilov2017decoupled), natural gradient descent (NGD) (amari1998natural), Gauss-Newton, K-FAC (martens2015optimizing), Shampoo (gupta2018shampoo), AdaGrad (duchi2011adaptive), RMSProp (tieleman2012lecture), and Muon (liu2025muon). In the following, we will present each optimizer in detail, with at least one corollary that mathematically registers each optimizer as a special case of our framework. The derivation follows the same structure as the main manuscript, extensively using the results delivered in Theorems 1 and 3, and Lemma 4. This reverse engineering not only allows us to uncover hidden design principles of these optimizers, but also places them in a unified framework, suggesting a systematic way to derive their optimal hyperparameters.
@@ E.1 Gradient descent
Gradient descent is a special case of Corollary 5. We present the result for the completeness of the framework. Since there is no momentum hyperparameter associated with gradient descent, the theorem only shows the connection between the budget (the trust region size) and the learning rate. In other words, the budget $B$ can be thought of as a theory-friendly notation for the learning rate $\eta$. In practical scenarios, we usually want to increase this value as much as possible to achieve the best performance and speed. Comparing this with Corollary 5, we get an insight of how the momentum hyperparameter is introduced to the stateless optimal optimizer, as well as how the backbone structure of the proofs for the optimal optimizers in this paper is constructed.
Corollary 11 (Instantaneously optimal GD).
@@
By Lemma 4, the optimal learning rate corresponds to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{SGD}},R\rangle_{{\mathcal{H}}}$ subject to the trust region constraint.
@@ E.2 Preconditioned gradient descent
Corollary 12 (Instantaneously optimal preconditioned gradient descent).
@@
By Lemma 4, the optimal preconditioned gradient descent parameters correspond to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{PGD}},R\rangle_{{\mathcal{H}}}$ subject to the elliptic trust region constraint.
@@
Keen readers will notice that the above proof is similar to the proof of optimal Adam in Corollary 6. Only the preconditioner $P$ is replaced by the sum of inverse per-parameter costs.
@@ E.3 Newton’s method
As we have discussed in the last section, we can derive the Newton’s method with momentum variable by changing the preconditioner $P$ to the inverse of the Hessian matrix $H^{-1}$ in the result of Corollary 12, assuming that the Hessian matrix $H\succ 0$ is positive definite and changing slowly compared to the gradient updates. We can also arbitrarily control the time-dependent memory components by changing the projection cone.
Corollary 13 (Instantaneously optimal Newton’s method (slow curvature change)).
@@
Proof of Corollary 13.
Setting $P=H^{-1}$ in the result of Corollary 12, we have the result. ∎
@@ E.4 Natural gradient descent
As we have discussed in the previous section, we can derive the natural gradient descent with momentum variable by changing the preconditioner $P$ to the inverse of the Fisher information matrix $F^{-1}$ in the result of Corollary 12, assuming that the Fisher information matrix $F\succ 0$ is positive definite and changing slowly compared to the gradient updates. We can also arbitrarily control the time-dependent memory components by changing the projection cone.
Corollary 14 (Instantaneously optimal natural gradient descent (slow curvature change)).
@@
Proof of Corollary 14.
Setting $P=F^{-1}$ in the result of Corollary 12, we have the result. ∎
@@ E.4.1 K-FAC
K-FAC (martens2015optimizing) is a block-diagonal Fisher-scaled optimizer that approximates the natural gradient descent by approximating the Fisher information matrix as a block-diagonal matrix. We assume that the Fisher information matrices are positive definite and changing slowly compared to the gradient updates, so we can approximate its inverse as a constant. Like the full NGD, we can derive K-FAC using our framework, extending the result of Corollary 12 to the block-diagonal case.
Corollary 15 (Instantaneously optimal K-FAC).
@@
Proof of Corollary 15.
Setting $P=\mathrm{bdiag}(F_{\ell}^{-1})$ in the preconditioned gradient descent result in Corollary 12, we have the result. ∎
@@ E.5 Shampoo
Shampoo (gupta2018shampoo) is a Kronecker-factored optimizer that approximates arbitrary color gradient descent (with or without momentum) by approximating the second moment matrix $P$ as a Kronecker product of mode-wise second moment matrices. Like all above, we can derive Shampoo using our framework, assuming that the second moment matrices are positive definite and changing slowly compared to the gradient updates.
Corollary 16 (Instantaneously optimal Shampoo).
@@
Proof of Corollary 16.
Setting $P=\bigotimes_{i}G_{i}^{-1/2}$ in the result of Corollary 12, we have the result. In addition, we have the impulse response $q_{k}=\eta(1-\beta)\beta^{k}\bigotimes_{i}G_{i}^{-1/2}$. By definition,
@@ E.6 Adaptive-moment family
The adaptive-moment family includes AdaGrad (duchi2011adaptive; agarwal2019efficient), RMSProp (tieleman2012lecture), and Adam (kingma2015adam). We have already derived Adam in Corollary 6 in the main text. This section completes the derivation of its family, which share the same proof structure. We assume that the second moment matrices change slower than the gradients, which is typical in practice.
@@ E.6.1 AdaGrad
Regarding AdaGrad (duchi2011adaptive; agarwal2019efficient), we have the expensive full-matrix version and the relatively cheap diagonal version.
Corollary 17 (Instantaneously optimal full-matrix AdaGrad with momentum).
@@
By plugging in $P=G$ in the result of Corollary 12, we have the exact same structure. This gives the desired results. ∎
Corollary 18 (Instantaneously optimal diagonal AdaGrad with momentum).
@@
By Lemma 4, the optimal AdaGrad parameters correspond to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{AdaGrad}},R\rangle_{{\mathcal{H}}_{D}}$ subject to the diagonal weighted trust region constraint.
@@ E.6.2 RMSProp
The RMSProp optimizer (tieleman2012lecture) can also be registered into our framework as a diagonal optimizer family, similarly to AdaGrad in Corollary 18. This is interesting because RMSProp does not use an explicit first-order moment, unlike many of the optimizers we have studied.
Corollary 19 (Instantaneously optimal RMSProp-style scaling).
@@ E.7 Muon
The Muon optimizer (liu2025muon) is an approximated preconditioned SGD with momentum, which is recently proposed for training large language models, designed to replace the former widely used alternative: AdamW (loshchilov2017decoupled). It combines momentum with an orthogonalization step that approximates the orthogonal polar factor of the momentum matrix, effectively performing steepest descent under a spectral-norm trust region.
@@
Corollary 20 (Instantaneously optimal Muon).
@@ F.3 Four families of optimal dynamic optimizers (Supplement to Section 3)
Instantiating the construction from Theorem 3 on each of these families, we obtain the closed-form optimal dynamic optimizer $Q^{\star}$ and the corresponding optimal learning power $P^{*}(R)$.
Corollary 21 (Closed-form solutions for dynamic trust regions).
@@
We apply Theorem 3 to each dynamic feasible set.
@@
The proof is similar to the proof of Corollary 2 of the main manuscript, which is provided in Appendix G. These analytic solutions reveal how the characteristics of different types of optimal dynamic optimizers $Q^{\star}$ are induced by controlling the feasible set $\mathcal{Q}$. Interpretation of the solutions is the same as the interpretation of the corresponding stateless solutions in Section 2.
@@ Appendix G Proofs omitted from the main text
G.1 Proof of Theorem 1
Proof of Theorem 1.
@@ G.1 Proof of Theorem 1
G.2 Proof of Corollary 2
Proof of Corollary 2.
We apply Theorem 1 to each type of trust regions. Let $\Sigma=U\operatorname{diag}(\sigma_{1}\geq\cdots\geq\sigma_{d})U^{\top}$ be the eigendecomposition of the moment matrix.
@@ G.2 Proof of Corollary 2
G.3 Proof of Theorem 3
Proof of Theorem 3.
We establish each claim in turn. The proof is largely similar to Theorem 1 due to the similarity of the problem settings, except for the type of inner product being used (Frobeinus vs. Hibert). Here, we give the full proof for completeness.
@@ G.3 Proof of Theorem 3
G.4 Proof of Lemma 4
Proof of Lemma 4.
The proof follows from standard convex optimization theory (boyd2004convex), specifically the KKT conditions for linear maximization and the characterization of metric projections.
@@ G.4 Proof of Lemma 4
G.5 Proof of Corollary 5
Proof of Corollary 5.
@@ G.5 Proof of Corollary 5
The 1-pole momentum family forms a feasible submanifold within the trust region constraint ${\mathcal{Q}}=\{Q:\|Q\|_{{\mathcal{H}}}\leq\sqrt{B}\}$ and the causality cone $\mathcal{C}$}. By Lemma 4, the optimal momentum parameters correspond to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{SGD+M}},R\rangle_{{\mathcal{H}}}$ subject to the trust region constraint.
@@
G.6 Proof of Corollary 6
Proof of Corollary 6.
@@ G.6 Proof of Corollary 6
The diagonal 1-pole Adam family forms a feasible submanifold within the diagonal trust region constraint ${\mathcal{Q}}_{D}$ and the causality cone $\mathcal{C}$}. By Lemma 4, the optimal Adam parameters correspond to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{Adam}},R\rangle_{{\mathcal{H}}}$ subject to the diagonal trust region constraint.
@@
G.7 Proof of Proposition 7
Proof of Proposition 7.
@@ G.7 Proof of Proposition 7
G.8 Proof of Lemma 8
For the proof of Lemma 8, we use the following lemma that formalizes the concept of alignment between the optimal optimizer and the gradient statistics as the sharing of common eigenstructures. The results are trivial from Theorem 1 but we state them for completeness.
Lemma 22 (Greedy optimizers are gradient-aligned).
@@ G.8 Proof of Lemma 8
Proof of Lemma 22.
We prove each case by applying the optimality conditions found in Theorem 1.
@@
Now we prove the commutativity of the optimal optimizer and the gradient moment in Lemma 8 of the main text.
Proof of Lemma 8.
We prove commutativity for each family from Lemma 22.
Case (i): Frobenius ball. From Lemma 22, $Q_{\textnormal{F}}^{\star}=\sqrt{B}\,R/\|R\|_{\textnormal{F}}$. Since $Q_{\textnormal{F}}^{\star}$ is a scalar multiple of $R$, they trivially commute:
@@
G.9 Proof of Theorem 9
Proof of Theorem 9.
@@ G.9 Proof of Theorem 9
Step 5 — Commutation refinement for optimal optimizers. For the Frobenius, spectral, data-metric, and diagonal families from Corollary 2, the optimal optimizer $Q^{\star}$ satisfies a key commutativity property with the Gram matrix $G=J^{\top}J$:
@@
This commutativity $Q^{\star},G]=0$ implies $K_{Q^{\star}},K_{I}]=0$ where $K_{I}=JJ^{\top}$ is the standard NTK (jacot2018neural). Hence $K_{Q^{\star}}$ and $K_{I}$ are simultaneously diagonalizable, and the kernel quantities separate mode-wise, enabling the water-filling closed forms under the trust region. ∎

### v2

@@ C.2 Extended experimental results
Method GSM8K acc. (%) Train loss
$\beta_{1}=0.5$ 52.57 $\pm$ 1.10 0.2080 $\pm$ 0.0004
$\beta_{1}=0.8$ 52.31 $\pm$ 1.00 0.2080 $\pm$ 0.0004
$\beta_{1}=0.9$ 51.76 $\pm$ 0.99 0.2081 $\pm$ 0.0004
$\beta_{1}=0.95$ 51.12 $\pm$ 0.77 0.2085 $\pm$ 0.0004
$\beta_{1}=0.98$ 51.25 $\pm$ 0.38 0.2093 $\pm$ 0.0005
$\beta_{1}=0.99$ 50.97 $\pm$ 0.68 0.2103 $\pm$ 0.0004
Ours 52.77 $\pm$ 0.93 0.2084 $\pm$ 0.0003
@@ Appendix D Automating Validation-Aware Optimizer Tuning
Proposition D.1 (Validation optimality in power).
@@
Figure 5: Demonstration of effectiveness of validation-aware design of gradient-based optimizers. The validation-aware optimizers achieve the highest test accuracy among all optimizers. The SGD+M optimizer is trained on the CIFAR-100 dataset (Krizhevsky, 2009) with ResNet-18 (He et al., 2016).
@@
Although, from Proposition D.1 and from our experiment, we know that validation-aware optimal optimizer gives instantaneously favorable choice over the training-only optimizer, a deeper analysis is required to understand and measure the exact effect of validation-awareness. We leave this as future work and resort to empirical evaluation for now. This concludes our theoretical and empirical demonstration of automatic, validation-aware design of gradient-based optimizers.
@@ Appendix E Reverse Engineering Common Optimizers
In the main manuscript, we have provided the optimal forms of two of the most widely used optimizers: SGD with momentum and Adam (Kingma and Ba, 2015), and postponed the extension to other optimizers to this appendix, leaving only Table 7 with the full list of optimizers in the main text. This section completes the reverse engineering of various optimizers under our greedy paradigm.
Our results include, but are not limited to, the following optimizers: SGD, Adam/AdamW (Kingma and Ba, 2015; Loshchilov and Hutter, 2019), natural gradient descent (NGD) (Amari, 1998), Gauss-Newton, K-FAC (Martens and Grosse, 2015), Shampoo (Gupta et al., 2018), AdaGrad (Duchi et al., 2011), RMSProp (Tieleman and Hinton, 2012), and Muon (Liu and others, 2025). In the following, we will present each optimizer in detail, with at least one corollary that mathematically registers each optimizer as a special case of our framework. The derivation follows the same structure as the main manuscript, extensively using the results delivered in Theorems 2.1 and 3.1, and Lemma 3.2. This reverse engineering not only allows us to uncover hidden design principles of these optimizers, but also places them in a unified framework, suggesting a systematic way to derive their optimal hyperparameters.
@@ E.1 Gradient descent
Gradient descent is a special case of Corollary 3.3. We present the result for the completeness of the framework. Since there is no momentum hyperparameter associated with gradient descent, the theorem only shows the connection between the budget (the trust region size) and the learning rate. In other words, the budget $B$ can be thought of as a theory-friendly notation for the learning rate $\eta$. In practical scenarios, we usually want to increase this value as much as possible to achieve the best performance and speed. Comparing this with Corollary 3.3, we get an insight of how the momentum hyperparameter is introduced to the stateless optimal optimizer, as well as how the backbone structure of the proofs for the optimal optimizers in this paper is constructed.
Corollary E.1 (Instantaneously optimal GD).
@@
By Lemma 3.2, the optimal learning rate corresponds to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{SGD}},R\rangle_{{\mathcal{H}}}$ subject to the trust region constraint.
@@ E.2 Preconditioned gradient descent
Corollary E.2 (Instantaneously optimal preconditioned gradient descent).
@@
By Lemma 3.2, the optimal preconditioned gradient descent parameters correspond to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{PGD}},R\rangle_{{\mathcal{H}}}$ subject to the elliptic trust region constraint.
@@
Keen readers will notice that the above proof is similar to the proof of optimal Adam in Corollary 3.4. Only the preconditioner $P$ is replaced by the sum of inverse per-parameter costs.
@@ E.3 Newton’s method
As we have discussed in the last section, we can derive the Newton’s method with momentum variable by changing the preconditioner $P$ to the inverse of the Hessian matrix $H^{-1}$ in the result of Corollary E.2, assuming that the Hessian matrix $H\succ 0$ is positive definite and changing slowly compared to the gradient updates. We can also arbitrarily control the time-dependent memory components by changing the projection cone.
Corollary E.3 (Instantaneously optimal Newton’s method (slow curvature change)).
@@
Proof of Corollary E.3.
Setting $P=H^{-1}$ in the result of Corollary E.2, we have the result. ∎
@@ E.4 Natural gradient descent
As we have discussed in the previous section, we can derive the natural gradient descent with momentum variable by changing the preconditioner $P$ to the inverse of the Fisher information matrix $F^{-1}$ in the result of Corollary E.2, assuming that the Fisher information matrix $F\succ 0$ is positive definite and changing slowly compared to the gradient updates. We can also arbitrarily control the time-dependent memory components by changing the projection cone.
Corollary E.4 (Instantaneously optimal natural gradient descent (slow curvature change)).
@@
Proof of Corollary E.4.
Setting $P=F^{-1}$ in the result of Corollary E.2, we have the result. ∎
@@ E.4.1 K-FAC
K-FAC (Martens and Grosse, 2015) is a block-diagonal Fisher-scaled optimizer that approximates the natural gradient descent by approximating the Fisher information matrix as a block-diagonal matrix. We assume that the Fisher information matrices are positive definite and changing slowly compared to the gradient updates, so we can approximate its inverse as a constant. Like the full NGD, we can derive K-FAC using our framework, extending the result of Corollary E.2 to the block-diagonal case.
Corollary E.5 (Instantaneously optimal K-FAC).
@@
Proof of Corollary E.5.
Setting $P=\mathrm{bdiag}(F_{\ell}^{-1})$ in the preconditioned gradient descent result in Corollary E.2, we have the result. ∎
@@ E.5 Shampoo
Shampoo (Gupta et al., 2018) is a Kronecker-factored optimizer that approximates arbitrary color gradient descent (with or without momentum) by approximating the second moment matrix $P$ as a Kronecker product of mode-wise second moment matrices. Like all above, we can derive Shampoo using our framework, assuming that the second moment matrices are positive definite and changing slowly compared to the gradient updates.
Corollary E.6 (Instantaneously optimal Shampoo).
@@
Proof of Corollary E.6.
Setting $P=\bigotimes_{i}G_{i}^{-1/2}$ in the result of Corollary E.2, we have the result. In addition, we have the impulse response $q_{k}=\eta(1-\beta)\beta^{k}\bigotimes_{i}G_{i}^{-1/2}$. By definition,
@@ E.6 Adaptive-moment family
The adaptive-moment family includes AdaGrad (Duchi et al., 2011; Agarwal et al., 2019), RMSProp (Tieleman and Hinton, 2012), and Adam (Kingma and Ba, 2015). We have already derived Adam in Corollary 3.4 in the main text. This section completes the derivation of its family, which share the same proof structure. We assume that the second moment matrices change slower than the gradients, which is typical in practice.
@@ E.6.1 AdaGrad
Regarding AdaGrad (Duchi et al., 2011; Agarwal et al., 2019), we have the expensive full-matrix version and the relatively cheap diagonal version.
Corollary E.7 (Instantaneously optimal full-matrix AdaGrad with momentum).
@@
By plugging in $P=G$ in the result of Corollary E.2, we have the exact same structure. This gives the desired results. ∎
Corollary E.8 (Instantaneously optimal diagonal AdaGrad with momentum).
@@
By Lemma 3.2, the optimal AdaGrad parameters correspond to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{AdaGrad}},R\rangle_{{\mathcal{H}}_{D}}$ subject to the diagonal weighted trust region constraint.
@@ E.6.2 RMSProp
The RMSProp optimizer (Tieleman and Hinton, 2012) can also be registered into our framework as a diagonal optimizer family, similarly to AdaGrad in Corollary E.8. This is interesting because RMSProp does not use an explicit first-order moment, unlike many of the optimizers we have studied.
Corollary E.9 (Instantaneously optimal RMSProp-style scaling).
@@ E.7 Muon
The Muon optimizer (Liu and others, 2025) is an approximated preconditioned SGD with momentum, which is recently proposed for training large language models, designed to replace the former widely used alternative: AdamW (Loshchilov and Hutter, 2019). It combines momentum with an orthogonalization step that approximates the orthogonal polar factor of the momentum matrix, effectively performing steepest descent under a spectral-norm trust region.
@@
Corollary E.10 (Instantaneously optimal Muon).
@@ F.3 Four families of optimal dynamic optimizers (Supplement to Section 3)
Instantiating the construction from Theorem 3.1 on each of these families, we obtain the closed-form optimal dynamic optimizer $Q^{\star}$ and the corresponding optimal learning power $P^{*}(R)$.
Corollary F.1 (Closed-form solutions for dynamic trust regions).
@@
We apply Theorem 3.1 to each dynamic feasible set.
@@
The proof is similar to the proof of Corollary 2.2 of the main manuscript, which is provided in Appendix G. These analytic solutions reveal how the characteristics of different types of optimal dynamic optimizers $Q^{\star}$ are induced by controlling the feasible set $\mathcal{Q}$. Interpretation of the solutions is the same as the interpretation of the corresponding stateless solutions in Section 2.
@@ Appendix G Proofs Omitted from the Main Text
G.1 Proof of Theorem 2.1
Proof of Theorem 2.1.
@@
G.2 Proof of Corollary 2.2
Proof of Corollary 2.2.
We apply Theorem 2.1 to each type of trust regions. Let $\Sigma=U\operatorname{diag}(\sigma_{1}\geq\cdots\geq\sigma_{d})U^{\top}$ be the eigendecomposition of the moment matrix.
@@
G.3 Proof of Theorem 3.1
Proof of Theorem 3.1.
We establish each claim in turn. The proof is largely similar to Theorem 2.1 due to the similarity of the problem settings, except for the type of inner product being used (Frobeinus vs. Hibert). Here, we give the full proof for completeness.
@@
G.4 Proof of Lemma 3.2
Proof of Lemma 3.2.
The proof follows from standard convex optimization theory (Boyd and Vandenberghe, 2004), specifically the KKT conditions for linear maximization and the characterization of metric projections.
@@
G.5 Proof of Corollary 3.3
Proof of Corollary 3.3.
@@
The 1-pole momentum family forms a feasible submanifold within the trust region constraint ${\mathcal{Q}}=\{Q:\|Q\|_{{\mathcal{H}}}\leq\sqrt{B}\}$ and the causality cone $\mathcal{C}$}. By Lemma 3.2, the optimal momentum parameters correspond to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{SGD+M}},R\rangle_{{\mathcal{H}}}$ subject to the trust region constraint.
@@
G.6 Proof of Corollary 3.4
Proof of Corollary 3.4.
@@
The diagonal 1-pole Adam family forms a feasible submanifold within the diagonal trust region constraint ${\mathcal{Q}}_{D}$ and the causality cone $\mathcal{C}$}. By Lemma 3.2, the optimal Adam parameters correspond to the projection of the unconstrained optimizer $Q^{\star}$ onto this submanifold, which reduces to maximizing the inner product $\langle Q_{\text{Adam}},R\rangle_{{\mathcal{H}}}$ subject to the diagonal trust region constraint.
@@
G.7 Proof of Proposition 5.1
Proof of Proposition 5.1.
@@
G.8 Proof of Lemma 5.2
For the proof of Lemma 5.2, we use the following lemma that formalizes the concept of alignment between the optimal optimizer and the gradient statistics as the sharing of common eigenstructures. The results are trivial from Theorem 2.1 but we state them for completeness.
Lemma G.1 (Greedy optimizers are gradient-aligned).
@@
Proof of Lemma G.1.
We prove each case by applying the optimality conditions found in Theorem 2.1.
@@
Now we prove the commutativity of the optimal optimizer and the gradient moment in Lemma 5.2 of the main text.
Proof of Lemma 5.2.
We prove commutativity for each family from Lemma G.1.
Case (i): Frobenius ball. From Lemma G.1, $Q_{\textnormal{F}}^{\star}=\sqrt{B}\,R/\|R\|_{\textnormal{F}}$. Since $Q_{\textnormal{F}}^{\star}$ is a scalar multiple of $R$, they trivially commute:
@@
G.9 Proof of Theorem 5.3
Proof of Theorem 5.3.
@@
Step 5 — Commutation refinement for optimal optimizers. For the Frobenius, spectral, data-metric, and diagonal families from Corollary 2.2, the optimal optimizer $Q^{\star}$ satisfies a key commutativity property with the Gram matrix $G=J^{\top}J$:
@@
This commutativity $Q^{\star},G]=0$ implies $K_{Q^{\star}},K_{I}]=0$ where $K_{I}=JJ^{\top}$ is the standard NTK (Jacot et al., 2018). Hence $K_{Q^{\star}}$ and $K_{I}$ are simultaneously diagonalizable, and the kernel quantities separate mode-wise, enabling the water-filling closed forms under the trust region. ∎